OpenStudy (anonymous):

Explain the difference between using the trigonometric ratios (sin, cos, tan) to solve for a missing angle in a right triangle versus using the reciprocal ratios (sec, csc, cot). You must use complete sentences and any evidence needed (such as an example) to prove your point of view. @phi @Compassionate

OpenStudy (anonymous):

@Hero @e.mccormick @karatechopper Could you help?

OpenStudy (phi):

Conceptually there is no difference. if you learn about ratios and equations like \[ \frac{a}{b} = \frac{c}{d} \] that is the same thing as \[ \frac{b}{a} = \frac{d}{c} \] and the same is true for (example) cosine versus secant: just "flip" both sides of the equation to switch between the two forms. \[ \cos x = \frac{adj}{hyp} \leftrightarrow \sec x = \frac{hyp}{adj}\] the only reason to prefer one ratio versus the other is to avoid a divide by zero.